Antuan Company set the following standard costs for one unit of its product.

Direct materials (5.0 Ibs. @ $5.00 per Ib.) $ 25.00
Direct labor (1.8 hrs. @ $12.00 per hr.) 21.60
Overhead (1.8 hrs. @ $18.50 per hr.) 33.30
Total standard cost $ 79.90


The predetermined overhead rate ($18.50 per direct labor hour) is based on an expected volume of 75% of the factory’s capacity of 20,000 units per month. Following are the company’s budgeted overhead costs per month at the 75% capacity level.

Overhead Budget (75% Capacity)
Variable overhead costs
Indirect materials $ 15,000
Indirect labor 75,000
Power

15,000

Repairs and maintenance 30,000
Total variable overhead costs $ 135,000
Fixed overhead costs
Depreciation—Building 25,000
Depreciation—Machinery 70,000
Taxes and insurance 18,000
Supervision 251,500
Total fixed overhead costs 364,500
Total overhead costs $ 499,500


The company incurred the following actual costs when it operated at 75% of capacity in October.

Direct materials (75,500 Ibs. @ $5.20 per lb.) $ 392,600
Direct labor (21,000 hrs. @ $12.20 per hr.) 256,200
Overhead costs
Indirect materials $ 41,000
Indirect labor 176,900
Power 17,250
Repairs and maintenance 34,500
Depreciation—Building 25,000
Depreciation—Machinery 94,500
Taxes and insurance 16,200
Supervision 251,500 656,850
Total costs $ 1,305,650

Required:
1&2. Prepare flexible overhead budgets for October showing the amounts of each variable and fixed cost at the 65%, 75%, and 85% capacity levels and classify all items listed in the fixed budget as variable or fixed.
Answer #1
Variable cost per unit Fixed cost 65% (13,000 units) 75% (15,000 units) 85% (17,000 units)
Variable overhead cost:
Indirect material ($15,000/15,000)=$1 $1 $13,000 $15,000 $17,000
Indirect labor (75,000/15,000)=$5 5 65,000 75,000 85,000
Power (15,000/15,000)=$1 1 13,000 15,000 17,000
Repairs and maintenance (30,000/15,000)=$2 2 26,000 30,000 34,000
Total variable overhead cost 9 117,000 135,000 153,000
Fixed overhead cost:
Depreciation-Building $25,000 25,000 25,000 25,000
Depreciation-Machinery 70,000 70,000 70,000 70,000
Taxes and insurance 18,000 18,000 18,000 18,000
Supervision 251,500 251,500 251,500 251,500
Total fixed overhead cost 364,500 364,500 364,500 364,500
Total overhead costs $481,500 $499,500 $517,500